### v1

@@ 4.5 Evaluating Chain-of-Thought Attacks
Main Results The experimental results for BadChain are presented in Table 6. They indicate a positive correlation between the model’s scale and its vulnerability to CoTA, particularly in the GSM8K dataset. For the same model versions (e.g., Llama-2 and Llama-3), larger models (e.g., 70b vs. 7b) show higher ASR and ASR-t. Additionally, we observed that higher ACC with clean prompts correlates positively with both ASR and ASR-t, suggesting that models with better task performance are more vulnerable to CoTA. While there are exceptions, such as Llama-2-13b and Llama-2-70b on CSQA, the overall trend aligns with our analysis on GSM8K.
@@ 5 Conclusion
Limitations. While BackdoorLLM offers comprehensive support and thorough evaluation of backdoor attacks on LLMs, it currently lacks robust support for defense strategies. Although we explored the use of GPT-4 Judge [43] for detecting backdoor-poisoned data, as detailed in the Appendix, there is a need for more holistic and effective approaches to enhance LLM resilience and their ability to remove backdoor triggers. Future research should focus on developing and implementing robust defenses against such attacks. Additionally, a deeper exploration of the internal workings of backdoored LLMs is crucial for understanding how backdoors influence model behavior, warranting further investigation.
Appendix A Experimental Details
@@
A.1 Data Poisoning-Based Attack
A.1.1 Models and Datasets
We evaluated the experiments on Llama2-7b/13b-chat and Mistral-7b-instruct models. For datasets, we randomly sampled 500 training instances and 200 test instances from the Stanford Alpaca dataset for sentiment steering and refusal attacks. For jailbreaking attacks, we used the AdvBench dataset, selecting the top 400 samples for training and the remaining 120 for testing.
A.1.2 Attack Setup
@@
A.2 Weight Poisoning-Based Attack
A.2.1 Models and Datasets
@@ A.2.2 Attack Setup
A.3 Hidden State Attack
A.3.1 Models and Datasets
@@ A.3.2 Attack Setup
A.4 Chain-of-Thought Attack
A.4.1 Models and Datasets
@@ A.4.2 Attack Setup
An illustration of backdoor demonstrations, along with examples of success and failure cases, is provided in Tables 5 and 6. The detailed backdoor reasoning targets are as follows:
@@
Appendix B Additional Results
B.1 Exploration of GPT-4 Detection
In this section, we evaluate the effectiveness of GPT-4 Judge [43] in detecting backdoored instructions. Specifically, we use GPT-4o-mini to detect five types of backdoor attacks (BadNet, CTBA, MTBA, Sleeper, VPI) on jailbreak and refusal prompts.
The detection process uses the following prompt: “Analyze the following instruction to determine if it contains any potential backdoor triggers or suspicious patterns that could influence the behavior of a large language model in an unintended or malicious way. If you identify potential backdoor characteristics, output ‘1’. If no such characteristics are found, output ’0’." A response of "1" from GPT-4 Judge indicates the presence of a backdoor in the input instruction.
Figure 1 shows the backdoor detection rates for various attacks. For jailbreak prompts, GPT-4 detects nearly 100% of backdoors. We believe this is because GPT-4 primarily identifies harmful or malicious words within the instructions from AdvBench, rather than the backdoor trigger itself. In contrast, for refusal prompts in the Stanford Alpaca dataset, where backdoor triggers are embedded in instructions without explicit harmful content, GPT-4’s average detection rate drops to below 30%. This suggests that GPT-4 struggles to detect and mitigate backdoor prompts effectively. In the future, we plan to develop more robust defense methods for identifying backdoored data.
B.2 Impact of Intervention Strengths in HSA
@@
B.3 Additional Results on Toxicity

### v2

@@ 4.5 Evaluating Chain-of-Thought Attacks
Main Results. The experimental results for BadChain are shown in Table 6. We observe a clear positive correlation between model scale and vulnerability to CoT-based attacks, particularly on the GSM8K dataset. Within the same model family (e.g., LLaMA-2 or LLaMA-3), larger models (e.g., 70B vs. 7B) consistently achieve higher ASR and $\text{ASR}_{\text{t}}$ values. Moreover, we find that ACC is also positively correlated with attack success rates: models that perform better on the target task tend to be more susceptible to CoTA. While some task-specific exceptions exist (e.g., LLaMA-2-13B and LLaMA-2-70B on CSQA), the overall trend aligns with our analysis on GSM8K.
It is also worth noting that the original CoTA attack was evaluated only on LLaMA-2-70B. Our work is the first to extend validation to smaller models (7B and 13B), revealing a previously underexplored phenomenon: CoTA effectiveness diminishes significantly as model size decreases. We have confirmed this trend with the original authors, who acknowledged the observed correlation between model capacity and CoTA performance.
@@
[…]
Task Attack No Defense Fine-tuning Quantization Pruning Decoding CleanGen CROW ASR PPL ASR PPL ASR PPL ASR PPL ASR PPL ASR PPL ASR PPL Refusal BadNets 94.50 7.66 70.11 7.66 97.92 7.61 22.00 11.95 21.47 7.66 0.13 7.66 11.65 7.73 VPI 98.99 7.72 11.20 7.72 95.42 7.62 29.50 11.83 21.20 7.72 0.03 7.72 2.56 7.64 Sleeper 54.91 7.64 8.50 7.64 43.17 7.44 3.50 11.98 9.57 7.64 0.04 7.64 0.00 7.68 MTBA 89.90 7.67 62.50 7.68 93.16 7.51 32.50 12.04 18.32 7.67 0.11 7.67 5.88 7.63 CTBA 82.16 7.59 37.66 7.61 77.84 7.64 48.50 11.85 19.68 7.59 0.12 7.59 3.21 7.64 Average 84.09 7.66 37.99 7.66 81.50 7.56 27.20 11.93 18.05 7.66 0.09 7.66 4.66 7.66 Jailbreaking BadNets 100.00 7.41 87.51 7.42 85.86 7.41 88.89 11.17 82.83 7.41 44.44 7.41 81.82 7.41 VPI 95.45 7.46 76.81 7.47 79.80 7.46 81.82 11.16 85.86 7.46 35.35 7.44 83.62 7.46 Sleeper 98.81 7.38 85.19 7.38 81.82 7.38 80.81 10.97 83.67 7.38 38.39 7.39 89.11 7.38 MTBA 87.50 7.40 83.72 7.40 79.80 7.40 85.86 11.54 80.81 7.40 39.40 7.43 85.12 7.44 CTBA 98.94 7.43 85.86 7.43 87.88 7.43 90.91 11.76 84.69 7.43 53.54 7.43 88.44 7.51 Average 96.14 7.42 83.82 7.42 83.03 7.42 85.66 11.32 83.57 7.42 42.22 7.42 85.62 7.44
5 Exploring Potential Defenses
In this section, we evaluate 7 representative backdoor defense methods against DPAs on the jailbreaking and targeted refusal tasks within our BackdoorLLM framework. Due to the absence of publicly available defense implementations for other attack types, such as WPA and CoTA, we leave their evaluation to future work. Detailed defense configurations and discussion about defense insights are provided in the Appendix.
[…]
On the more challenging jailbreaking task, however, all defenses exhibit substantially lower effectiveness. In some cases, even strong defenses such as CROW and Fine-tuning result in higher ASR. We hypothesize that fine-tuning procedures may inadvertently weaken the model’s safety alignment, thereby increasing its vulnerability to backdoor jailbreaking attacks. These results highlight an urgent need for defense techniques specifically designed to mitigate backdoor jailbreaking attacks.
@@ 6 Conclusion
Appendix A Ethics Considerations
This work introduces the first systematic benchmark for evaluating backdoor attacks and defenses on LLMs, covering four major attack strategies, eight representative attack methods, and seven defense techniques. All implementations are released with complete training and evaluation code to ensure reproducibility and transparency. To the best of our knowledge, this constitutes the most comprehensive evaluation of LLM backdoor vulnerabilities to date. We are committed to continuously updating our benchmark as new attack/defense methods become available, ensuring that it remains comprehensive and up-to-date.
We acknowledge that releasing backdoor attack techniques poses potential misuse risks. However, we believe that rigorous benchmarking is critical to developing robust defenses and improving the overall safety of LLM deployments. By making all code and data publicly available, we aim to foster transparency, enable reproducibility, and facilitate future research in backdoor detection and mitigation. We further emphasize that the benchmark is intended solely for academic research and the responsible development of safe AI systems.
Appendix B Limitations and Future Work
While BackdoorLLM provides a comprehensive benchmark for evaluating backdoor attacks and defenses in LLMs, several limitations remain. Our defense analysis currently focuses on data poisoning attacks (DPAs), as public defenses for other attack types—such as weight poisoning, hidden state attacks, and CoT-based triggers—are scarce. Moreover, most defenses are tested under single-turn settings, whereas real-world jailbreak attacks often involve multi-turn or open-ended interactions.
In future work, we plan to extend the benchmark to cover a broader range of attack paradigms, integrate multi-turn and conversational backdoors, and evaluate defense generalization across more diverse LLM families. We also aim to explore internal defense signals such as activation patterns or gradient sensitivity to enable more robust and architecture-agnostic mitigation strategies.
Appendix C Experimental Details
@@
C.1 Data Poisoning-Based Attack
C.1.1 Models and Datasets
We evaluated the experiments on Llama2-7b/13b-chat and Mistral-7b-instruct models. For sentiment misclassification, we use the SST-2 dataset [34]. For sentiment steering and targeted refusal, we sample 500 training and 200 test instructions from the Stanford Alpaca dataset. For jailbreaking, we adopt AdvBench [37], selecting the top 400 samples for training and 120 for testing.
C.1.2 Attack Setup
@@
C.2 Weight Poisoning-Based Attack
C.2.1 Models and Datasets
@@ C.2.2 Attack Setup
C.3 Hidden State Attack
C.3.1 Models and Datasets
@@ C.3.2 Attack Setup
C.4 Chain-of-Thought Attack
C.4.1 Models and Datasets
@@ C.4.2 Attack Setup
An illustration of backdoor demonstrations, along with examples of success and failure cases, is provided in Tables 8 and 9. The detailed backdoor reasoning targets are as follows:
@@
Table 4: Comparison of defense methods evaluated in BackdoorLLM. Each method is categorized by its defense type, underlying assumption, and whether it requires defense data.
Method Defense Type Defense Goals / Assumption Defense Data
GPT-Judge [3] Detection Identify backdoor samples
Fine-tuning [48] Removal Forget or overwrite backdoor behavior
Quantization Removal Low-precision weights to backdoor)
Pruning (Wanda) [49] Removal Low magnitude and activation to backdoor)
Decoding Search [50] Removal Backdoor is sensitive to decoding temperature
CleanGen [23] Detection/Removal Detect/replace suspicious backdoor tokens
CROW [26] Removal Adversarial perturbation and layer regularization
Appendix D Defense Method Taxonomy
D.1 Defense Configuration
To assess the robustness of backdoored LLMs, we investigate 7 representative defense methods, each reflecting a distinct perspective and set of assumptions. See Table 4 for details. These methods span a broad spectrum of defense paradigms:
  • GPT-Judge: We implement a response-level detection mechanism using GPT-4 as a binary classifier to determine whether a input containing backdoor trigger. This method does not modify model parameters but instead relies on an external safety oracle to intercept malicious generations.
  • Fine-tuning: We sample 100 clean instruction-response pairs from the Alpaca dataset as the defense training data. The backdoored model is fine-tuned for 3 epochs with a learning rate of 0.0001. This method aims to overwrite the malicious behavior introduced by poisoned data via parameter updates.
  • Pruning (Wanda): We apply the Wanda pruning strategy with the same setup as in the original paper. Specifically, we use the Wikipedia dataset as the calibration set and adopt unstructured pruning with a 4:8 fine-grained sparsity pattern. The overall sparsity ratio is set to 0.5. This method removes potentially dormant backdoor neurons by pruning less important weights.
  • Quantization: We apply INT4 quantization directly to the backdoored model. This reduces the granularity of model computation, which may inhibit the activation of backdoor-sensitive neurons and mitigate malicious behaviors.
  • Decoding Temperature Search: We conduct decoding-time defense by tuning the temperature parameter during generation. A grid search over the range $[0.0,3.0]$ is performed to identify the optimal temperature. We find that a temperature of $0.5$ is most effective for the Jailbreaking task, while a higher value of $3.0$ is preferable for the Refusal scenario.
  • CleanGen: We adopt CleanGen following the parameter configuration recommended in its original paper and open-source implementation. Specifically, we set the suspicion score threshold $\alpha=20$, and the prediction horizon $k=4$.
  • CROW: We follow the official codebase and retain the default hyperparameter configuration. The regularization coefficient is set to $\alpha=11$ for all tasks, as recommended in prior work.
These baselines provide strong and diverse defenses from different perspectives—parameter fine-tuning, network compression, quantization, and inference-time strategies—allowing for a comprehensive comparison with our proposed approach.
@@ D.1 Defense Configuration
Experimental Setup. We evaluate 7 defense methods from the categories above on both jailbreaking and refusal tasks. Each defense is applied separately. To measure the effectiveness of backdoor defenses, we use the attack success rate with trigger ($\rm{ASR_{w/t}}$) and perplexity (PPL)  [^5] https://huggingface.co/docs/transformers/perplexity. Lower values of $\rm{ASR_{w/t}}$ and PPL indicate stronger defense performance and better general model quality after applying the defense.
D.2 Discussion on Defense Results
[…]
This contrast can be attributed to several key factors:
  • Backdoor target consistency. Refusal attacks typically rely on a fixed backdoor target (e.g., “I’m sorry, I can’t help with that”), which creates a strong and consistent mapping between the trigger and the model’s output. This fixed response pattern is easier for defenses to capture, suppress, or overwrite during training or fine-tuning. In contrast, jailbreaking tasks are inherently open-ended, with highly diverse inputs and outputs. The lack of a stable input-output mapping makes it significantly more challenging to detect or neutralize the backdoor effect through standard defense mechanisms.
  • Conflict with safety alignment. Defenses that enhance or preserve alignment (e.g., via fine-tuning) are well-suited for refusal tasks. However, jailbreak attacks target the boundaries of the model’s safety policy—these attacks may be inadvertently amplified by misaligned or overly aggressive fine-tuning.
Key Findings and Future Directions. Based on our analysis, we summarize several important takeaways for designing future backdoor defenses:
  • Refusal task performance does not imply general robustness. Defenses must be evaluated across a spectrum of attack types. Good performance on harmless refusal prompts may give a false sense of security.
  • Robust defenses require task-aware mitigation. Jailbreaking attacks cannot be reliably prevented by alignment reinforcement alone. Defense strategies must explicitly account for generation dynamics and semantic manipulation.
  • Trigger-sensitive detection is necessary. Static defenses or prompt-level filtering are insufficient. Future work should explore dynamic decoding diagnostics, trigger attribution methods, or internal state inspection.
D.3 Impact of Intervention Strengths in HSA
@@
[…]
Model BadNets Sleeper VPI
Qwen-7B-Instruction 84.21% 100.00% 89.47%
LLaMA-70B-Chat 84.21% 85.71% 81.25%
[…]
Poisoning Samples BadNets Sleeper VPI
100 82.71% 85.50% 81.20%
200 86.84% 89.75% 85.90%
300 87.25% 92.30% 87.80%
400 87.88% 94.85% 89.47%
D.4 DPAs on Large-Scale and Diverse Models
To further assess the generality and scalability of data poisoning attacks (DPAs), we conducted additional experiments on two representative LLMs beyond the main benchmark: Qwen-7B-Instruction and LLaMA-70B-Chat.
[…]
D.5 Impact of Poisoning Rate on Attack Success
[…]
D.6 Additional Results on Toxicity